OpenStudy (anonymous):

What is the rate if the principal is $550, the time is 3 years, and the interest is $57.75 and how?

OpenStudy (lataesiaaa):

i believe its dookey.

OpenStudy (anonymous):

haha very funny

OpenStudy (mathmusician):

I just answered your question @lataesiaaa

OpenStudy (anonymous):

what???

OpenStudy (lataesiaaa):

its dookey with a side of eyeballs

OpenStudy (anonymous):

really this is not funny I need help

OpenStudy (loser66):

What is the formula to find out the interest ?

OpenStudy (anonymous):

I= p(principal) x r(rate) x t(time)

OpenStudy (loser66):

\(I = p *r* t\), then to find \(r\), you need divide both sides by \(p*t\) Now, what do you get?

OpenStudy (anonymous):

3.5%

OpenStudy (loser66):

\(r =\dfrac{I}{p*t}\) your \(I= 57.5\\p=550\\t=3\) nothing is hard, right?

OpenStudy (anonymous):

nope

OpenStudy (loser66):

hahaha.. ok, so it is hard to me

OpenStudy (anonymous):

yeah oh and it's 57.75 for interest

OpenStudy (loser66):

ok

OpenStudy (anonymous):

so I did it differently

OpenStudy (mathstudent55):

\(I = prt\) \(prt = I\) \(r = \dfrac{I}{pt} \) \(r = \dfrac{$57.75}{$550 \times 3} \) \(r = \dfrac{57.75}{1650} \) \(r = 0.035\) \(r = 3.5\%\)